# MIRA: Medical Time Series Foundation Model for Real-World Health Data

### v2

@@ Abstract
A unified foundation model for medical time series—pretrained on open access and ethics board-approved medical corpora—offers the potential to reduce annotation burdens, minimize model customization, and enable robust transfer across clinical institutions, modalities, and tasks, particularly in data-scarce or privacy-constrained environments. However, existing generalist time series foundation models struggle to handle medical time series data due to their inherent challenges, including irregular intervals, heterogeneous sampling rates, and frequent missing values. To address these challenges, we introduce MIRA, a unified foundation model specifically designed for medical time series forecasting. MIRA incorporates a Continuous-Time Rotary Positional Encoding that enables fine-grained modeling of variable time intervals, a frequency-specific mixture-of-experts layer that routes computation across latent frequency regimes to further promote temporal specialization, and a Continuous Dynamics Extrapolation Block based on Neural ODE that models the continuous trajectory of latent states, enabling accurate forecasting at arbitrary target timestamps. Pretrained on a large-scale and diverse medical corpus comprising over 454 billion time points collect from publicly available datasets, MIRA achieves reductions in forecasting errors by an average of 10% and 7% in out-of-distribution and in-distribution scenarios, respectively, when compared to other zero-shot and fine-tuned baselines. We also introduce a comprehensive benchmark spanning multiple downstream clinical tasks, establishing a foundation for future research in medical time series modeling.
@@ 1 Introduction
Medical time series data, including signals such as electrocardiograms (ECG) [1], electroencephalograms (EEG) [2], vital signs and laboratory measurements [3], are key to understanding the dynamic physiological states of patients over time [4, 5]. Continuously modeling these patient data trajectories supports clinical forecasting tasks such as predicting organ failure or treatment response [1, 6], enabling earlier decisions and better patient outcomes [7, 8]. However, effectively utilizing medical time series data in practice remains challenging, as patient populations, disease profiles, and clinical protocols can vary widely across regions and institutions. Moreover, such data is often collected in an irregular and asynchronous manner [9, 10], further complicating efforts to build generalizable models. These challenges are compounded by regulatory constraints such as GDPR [11, 12], which inhibit data sharing and result in institutions training task-specific models from scratch on local data, incurring significant resource overhead. This highlights the need for a shareable medical foundation model trained on validated, open medical time series datasets, that is able to learn generalizable patterns from large and varied datasets, reducing the need for extensive data annotation and custom model building, while also helping to share knowledge effectively across different clinical institutions, data types, and medical tasks [13, 14].
@@
To address this central question, we introduce MIRA—a Medical Foundation Model for Irregularity Adaptation. MIRA is specifically designed to address the challenges of medical time series: First, we propose a Continuous-Time Rotary Positional Encoding (CT-RoPE), which extends the standard RoPE [38] to handle time values using learnable frequency adjustments, allowing MIRA to accurately model the varying frequency intervals of irregularly sampled data. Second, to further improve its ability to adapt to different time features in medical signals, MIRA uses a frequency-specific mixture-of-experts (MoE) layer that learns to route frequency-related information through specialized components. Finally, MIRA integrates a Continuous Dynamics Extrapolation Block, using Neural Ordinary Differential Equations (Neural ODEs) [39]. This feature allows the model to predict patient health paths at any future time point, moving beyond the limits of fixed observation grids. MIRA is pretrained on a large curated collection of medical time series, comprised of over 454 billion time points from publicly available datasets. These datasets cover readings from intensive care units (ICUs), operating rooms, pediatric critical care, and long-term sleep and mental health monitoring. All data undergo preprocessing that standardizes time alignment, normalizes sampling frequencies, and maintains clinical realism. MIRA outperforms other time series foundation models with a similar number of activated parameters across various real-world benchmarks, achieving reductions in forecasting errors by an average of 10% and 7% in out-of-distribution and in-distribution scenarios, respectively.
@@ 3.2.1 Continuous-Time Rotary Positional Encoding
$\langle q_{m},k_{n}\rangle=x_{m}^{\top}(W^{Q})^{\top}R_{\Theta,m}^{\top}R_{\Theta,n}W^{K}x_{n}$ (4)
Since both $R_{\Theta,m}$ and $R_{\Theta,n}$ are rotation matrices constructed from trigonometric functions of the timestamp-derived angles, their product is itself a rotation matrix dependent only on the time difference:
$R_{\Theta,m}^{\top}R_{\Theta,n}=R_{\Theta,n-m}$ (5)
This implies that the positional interaction between two tokens is solely a function of their relative timestamp offset $(n-m)$. In CT-RoPE, this property generalizes to real-valued timestamps, enabling the attention mechanism to capture continuous-time relational structure while retaining the efficiency and structure of standard dot-product attention. In line with Chowdhery et al. [71], we remove biases except in QKV projections to enhance extrapolation. More detail can be found in Appendix C.
@@ 3.2.2 Frequency-Specific Mixture-of-Experts Block
$s_{i,t}$ $=\frac{\exp\left((\mathbf{W}_{i}^{l})^{\top}\mathbf{\bar{u}}^{l}_{t}\right)}{\sum_{j=1}^{N}\exp\left((\mathbf{W}_{j}^{l})^{\top}\mathbf{\bar{u}}^{l}_{t}\right)},\quad g_{i,t}$ =$=\begin{cases}s_{i,t},&\text{if }s_{i,t}\in\operatorname{TopK}(\{s_{j,t}\}_{j=1}^{N},K),\\ \\ 0,&\text{otherwise},\end{cases}$ (7)
@@ 3.3 Model Training
To support generalizable and clinically relevant time series modeling, we curate a large pretraining corpus spanning over 454 billion time points from various real-world healthcare settings. The dataset collection includes signals from ICUs a nd operating rooms, pediatric critical care, long-term sleep and mental health monitoring, and population-level epidemiological surveillance. All data are drawn from publicly available clinical datasets, including MIMIC-III [72], MIMIC-IV [73], PTB-XL [74], Sleep-EDF [75], and the WAVES Pediatric Waveform Database [76]. A full summary of included datasets and the pre-processing applied is provided in Appendix E.
@@ 4.1 Experiment Setup
Downstream Clinical Datasets. We evaluate MIRA on a diverse suite of real-world clinical and public health datasets, spanning multi-modal physiological signal analysis, critical care monitoring, ambulatory bio-signal tracking, epidemiological surveillance, pharmacological safety, and healthcare resource utilization. To systematically study temporal robustness, we group datasets into two categories: (1) inherently irregular datasets, which exhibit irregular sampling and natural missing values due to clinical or observational workflows. These are CinC 2012 [77] and CinC 2019 [78]. . Furthermore, we use (2) originally regular datasets, i.e.,MIT-BIH [79], Johns Hopkins COVID-19 Dataset [80] ,CDC Influenza Hospitalizations Admissions (CDC-IHA)  [^1] https://www.cdc.gov/flu-forecasting/data-vis/current-week.html, WHO fluent  [^2] https://www.who.int/tools/flunet, Heart Rate [81] and illness [82], for which we simulate irregularity by randomly masking 30% of time points. The full list of datasets and statistics is provided in Appendix G.
@@
Models Zero-shot Ours Full-shot Time Series Models
@@
Cinc 2012 ($10^{1}$) $5.963$ $5.891$ $5.764$ $5.732$ $5.312$ $5.120$ $5.987$ $5.985$ $6.247$ $6.246$ $6.997$ $6.995$ $7.498$ $7.497$
Cinc 2019 ($10^{1}$) $7.456$ $6.597$ $7.347$ $7.311$ $7.212$ $7.246$ $6.989$ $7.072$ $7.608$ $7.541$ $10.293$ $9.7724$ $9.591$ $8.322$
[…]
MIT-BIH $0.293$ $0.198$ $0.163$ $0.111$ $0.122$ $0.091$ $0.453$ $0.354$ $0.705$ $0.627$ $0.882$ $0.623$ $0.242$ $0.196$ $0.335$ $0.141$
CDC-IHA ($10^{1}$) $5.976$ $4.684$ $5.729$ $4.502$ $5.534$ $4.401$ $5.510$ $4.133$ $9.522$ $7.974$ $10.068$ $9.052$ $7.892$ $6.766$ $15.633$ $4.408$
@@ 4.2 Performance on Out-of-distribution Forecasting
Results. Detailed results of out-of-distribution performance are reported in Table 2. First, MIRA consistently achieves state-of-the-art performance, outperforming all general-domain foundation models and specialized time series baselines. Specifically, MIRAlarge achieves the best results on all datasets, improving RMSE by over 10% on average compared to the strongest baselines, confirming the advantage of scaling and medical pretraining. This advantage is particularly pronounced on clinical benchmarks such as MIT-BIH and CDC-IHA, where MIRAlarge achieves both the lowest RMSE and MAE. Second, domain-specific pretraining proves essential. All model variants—continue pretrained on medical corpora—consistently outperform models trained on general time series data. This demonstrates the benefit of leveraging medical-specific temporal structures and distributions during pretraining. Interestingly, even smaller MIRA models surpass larger general-domain models, suggesting that data relevance is more critical than model size alone. Third, MIRA achieves performance close to or even exceeding fine-tuned full-shot models in several cases. For instance, $MIRA_{large}$ outperform than full-shot baseline on Cinc 2012 and slight lower in JH Covid-19.
@@ 4.3 Performance on In-distribution Forecasting
Ours MIRAsmall $3.15$ $2.82$ $3.28$ $2.95$ $3.41$ $3.08$ $3.54$ $3.20$ $3.67$ $3.34$ $3.80$ $3.47$ $3.93$ $3.61$ $4.07$ $3.75$ $4.22$ $3.90$ 00 00
MIRAbase $2.98$ $2.68$ $3.11$ $2.81$ $3.25$ $2.94$ $3.38$ $3.07$ $3.51$ $3.20$ $3.64$ $3.34$ $3.78$ $3.48$ $3.93$ $3.63$ $4.10$ $3.79$ 00 $1$
MIRAlarge $2.85$ $2.56$ $2.95$ $2.65$ $3.08$ $2.78$ $3.21$ $2.90$ $3.35$ $3.04$ $3.49$ $3.18$ $3.63$ $3.31$ $3.99$ $3.86$ $3.97$ $3.64$ $8$ $8$
[…]
Time-MoElarge $3.20$ $2.91$ $3.33$ $3.04$ $3.47$ $3.17$ $3.60$ $3.29$ $3.73$ $3.41$ $3.86$ $3.54$ $3.99$ $3.67$ $4.13$ $3.81$ $4.28$ $3.96$ 00 00
Moirailarge $4.80$ $4.18$ $5.39$ $4.64$ $5.68$ $4.55$ $6.02$ $4.83$ $6.09$ $5.06$ $6.18$ $5.02$ $6.31$ $5.09$ $6.51$ $5.24$ $6.71$ $5.36$ 00 00
Moiraibase $4.84$ $4.15$ $4.93$ $4.10$ $5.20$ $4.23$ $5.64$ $4.53$ $5.78$ $4.61$ $5.80$ $4.60$ $5.95$ $4.97$ $5.95$ $4.73$ $6.06$ $4.87$ 00 00
Moiraismall $4.91$ $4.22$ $5.04$ $4.20$ $6.12$ $4.93$ $6.23$ $4.98$ $6.44$ $5.49$ $6.47$ $5.18$ $6.59$ $5.60$ $6.66$ $5.20$ $6.86$ $5.71$ 00 00
Chronoslarge $4.71$ $4.10$ $4.85$ $4.22$ $5.02$ $4.31$ $5.31$ $4.52$ $5.58$ $4.70$ $5.82$ $4.91$ $5.97$ $5.10$ $6.15$ $5.32$ $6.43$ $5.49$ 00 00
Chronosbase $5.02$ $4.40$ $5.21$ $4.51$ $5.43$ $4.68$ $5.62$ $4.81$ $5.85$ $4.93$ $6.02$ $5.12$ $6.21$ $5.31$ $6.40$ $5.49$ $6.67$ $5.62$ 00 00
Chronossmall $5.32$ $4.71$ $5.49$ $4.82$ $5.68$ $4.93$ $5.81$ $5.02$ $6.01$ $5.20$ $6.21$ $5.32$ $6.40$ $5.49$ $6.58$ $5.61$ $6.83$ $5.72$ 00 00
[…]
Time-MoEbase $3.05$ $2.78$ $3.18$ $2.91$ $3.32$ $3.04$ $3.45$ $3.17$ $3.58$ $3.30$ $3.71$ $3.43$ $3.85$ $3.57$ $3.99$ $3.71$ $4.15$ $3.86$ 00 00
Moirailarge $4.41$ $3.76$ $4.92$ $4.20$ $5.35$ $4.41$ $5.69$ $4.76$ $5.93$ $4.80$ $6.09$ $4.92$ $6.23$ $4.98$ $6.48$ $5.29$ $6.62$ $5.30$ 00 00
Moiraibase $4.55$ $3.92$ $4.78$ $4.07$ $5.18$ $4.26$ $5.49$ $4.46$ $5.78$ $4.64$ $5.70$ $4.51$ $5.87$ $4.97$ $5.92$ $4.75$ $5.96$ $4.71$ 00 00
Moiraismall $5.02$ $4.21$ $5.36$ $4.40$ $5.55$ $4.98$ $5.97$ $4.92$ $6.25$ $5.34$ $6.43$ $5.11$ $6.44$ $5.35$ $6.57$ $5.41$ $6.58$ $5.30$ 00 00
Chronoslarge $4.30$ $3.89$ $4.58$ $4.02$ $4.83$ $4.20$ $5.02$ $4.31$ $5.32$ $4.51$ $5.61$ $4.73$ $5.83$ $4.91$ $6.02$ $5.10$ $6.31$ $5.32$ 00 00
Chronosbase $4.61$ $4.10$ $4.83$ $4.20$ $5.02$ $4.31$ $5.32$ $4.51$ $5.58$ $4.70$ $5.82$ $4.91$ $5.97$ $5.10$ $6.15$ $5.32$ $6.43$ $5.49$ 00 00
Chronossmall $4.83$ $4.31$ $5.02$ $4.43$ $5.32$ $4.51$ $5.58$ $4.70$ $5.82$ $4.91$ $5.97$ $5.10$ $6.15$ $5.32$ $6.43$ $5.49$ $6.71$ $5.61$ 00 00
@@
Pretraining Scalability. Beyond downstream performance, we delve into the utilization of model capacity. We plot training curves in Figure 3. Compared with small size, the large version leads to totally 11% reduction in the converged training objective, showing significant performance promotion on in-distribution medical time series forecasting.

### v5

@@ Abstract
A foundation model for medical time series, pretrained on ethically approved clinical datasets, can substantially reduce annotation burdens, minimize the need for task-specific tuning, and promote reliable transferability across healthcare institutions, data modalities, and clinical tasks, especially in data-scarce or privacy-sensitive environments. However, existing generalist time series foundation models struggle to handle medical time series data due to their inherent challenges, including irregular intervals, heterogeneous sampling rates, and frequent missing values. To address these challenges, we introduce MIRA, a unified foundation model specifically designed for medical time series forecasting. MIRA incorporates a Continuous-Time Rotary Positional Encoding that enables fine-grained modeling of variable time intervals, a frequency-specific mixture-of-experts layer that routes computation across latent frequency regimes to further promote temporal specialization, and a Continuous Dynamics Extrapolation Block based on Neural ODE that models the continuous trajectory of latent states, enabling accurate forecasting at arbitrary target timestamps. Pretrained on a large-scale and diverse medical corpus comprising over 454 billion time points collect from publicly available datasets, MIRA achieves reductions in forecasting errors by an average of 8% and 6% in out-of-distribution and in-distribution scenarios, respectively, when compared to other zero-shot and fine-tuned baselines. We also introduce a comprehensive benchmark spanning multiple downstream clinical tasks, establishing a foundation for future research in medical time series modeling. Our code is available at MIRA.
@@ 1 Introduction
Medical time series data, including signals such as electrocardiograms (ECG) [1], electroencephalograms (EEG) [2], vital signs and laboratory measurements [3], are key to understanding the dynamic physiological states of patients over time [4, 5]. Continuously modeling these patient data trajectories supports clinical forecasting tasks such as predicting organ failure or treatment response [1, 6], enabling earlier decisions and better patient outcomes [7, 8]. However, effectively utilizing medical time series data in practice remains challenging, as patient populations, disease profiles, and clinical protocols can vary widely across regions and institutions [9]. Moreover, such data is often collected in an irregular and asynchronous manner [10, 11], further complicating efforts to build generalizable models. These challenges are amplified by regulatory frameworks such as GDPR [12, 13], which restrict cross-institutional data exchange and hinder the development of unified modeling paradigms, often resulting in redundant and resource-intensive local optimization efforts. This highlights the need for a foundation model trained on validated medical time series datasets, that is able to learn generalizable patterns from large and varied datasets, reducing the need for extensive data annotation and custom model building, while also helping to share knowledge effectively across different clinical institutions, data types, and medical tasks [14, 15].
@@
To address this central question, we introduce MIRA—a Medical Foundation Model for Irregularity Adaptation. MIRA is specifically designed to address the challenges of medical time series: First, we propose a Continuous-Time Rotary Positional Encoding (CT-RoPE), which extends the standard RoPE [39] to handle time values using learnable frequency adjustments, allowing MIRA to accurately model the varying frequency intervals of irregularly sampled data. Second, to further improve its ability to adapt to different time features in medical signals, MIRA uses a frequency-specific mixture-of-experts (MoE) layer that learns to route frequency-related information through specialized components. Finally, MIRA integrates a Continuous Dynamics Extrapolation Block, using Neural Ordinary Differential Equations (Neural ODEs) [40]. This feature allows the model to predict patient health paths at any future time point, moving beyond the limits of fixed observation grids. MIRA is pretrained on a large curated collection of medical time series, comprised of over 454 billion time points from publicly available datasets. These datasets cover readings from intensive care units (ICUs), operating rooms, pediatric critical care, and long-term sleep and mental health monitoring. All data undergo preprocessing that standardizes time alignment, normalizes sampling frequencies, and maintains clinical realism. MIRA outperforms other time series foundation models with a similar number of activated parameters across various real-world benchmarks, achieving reductions in forecasting errors by an average of 8% and 6% in out-of-distribution and in-distribution scenarios, respectively.
@@ 3.2.1 Continuous-Time Rotary Positional Encoding
$\langle q_{m},k_{n}\rangle\;=\;x_{m}^{\top}(W^{Q})^{\top}\big(R_{\Theta}(t_{m})\big)^{\top}R_{\Theta}(t_{n})\,W^{K}x_{n}.$ (4)
Since each $R_{\Theta,i}(t)$ is rotation matrices constructed from trigonometric functions of the timestamp-derived angles, their product is itself a rotation matrix dependent only on the time difference:
$\big(R_{\Theta}(t_{m})\big)^{\top}R_{\Theta}(t_{n})\;=\;R_{\Theta}(t_{n}-t_{m}),$ (5)
This implies that the positional interaction between two tokens is solely a function of their relative timestamp offset (nm)(n-m). In CT-RoPE, this property generalizes to real-valued timestamps, enabling the attention mechanism to capture continuous-time relational structure while retaining the efficiency and structure of standard dot-product attention. In line with Chowdhery et al. [72], we remove biases except in QKV projections to enhance extrapolation. More detail can be found in Appendix C.
@@ 3.2.2 Frequency-Specific Mixture-of-Experts Block
[…]
@@ 3.3 Model Training
To support generalizable and clinically relevant time series modeling, we curate a large pretraining corpus spanning over 454 billion time points from various real-world healthcare settings. The dataset collection includes signals from ICUs a nd operating rooms, pediatric critical care, long-term sleep and mental health monitoring, and population-level epidemiological surveillance. All data are drawn from publicly available clinical datasets, including MIMIC-III [73], MIMIC-IV [74], PTB-XL [75], Sleep-EDF [76], and the WAVES Pediatric Waveform Database [77]. To enable the model to acquire general knowledge, we resumed training from the Time-MoE checkpoint. A full summary of included datasets and the pre-processing applied is provided in Appendix E.
@@ 4.1 Experiment Setup
Downstream Clinical Datasets. We evaluate MIRA on a diverse suite of real-world clinical and public health datasets, spanning multi-modal physiological signal analysis, critical care monitoring, ambulatory bio-signal tracking, epidemiological surveillance, and healthcare resource utilization. To systematically study temporal robustness, we group datasets into two categories: (1) inherently irregular datasets, which exhibit irregular sampling and natural missing values due to clinical or observational workflows. These are CinC 2012 [78] and CinC 2019 [79]. Furthermore, we use (2) originally regular datasets, i.e.,MIT-BIH [80], Johns Hopkins COVID-19 Dataset [81] ,CDC Influenza Hospitalizations Admissions (CDC-IHA)  [^1] https://www.cdc.gov/flu-forecasting/data-vis/current-week.html, Heart Rate [82] and illness [83], for which we simulate irregularity by randomly masking 30% of time points. The full list of datasets and statistics is provided in Appendix G.
@@
Models
Zero-shot Ours Full-shot Time Series Models
@@
Cinc 2012 ($10^{1}$) $7.136$ $6.984$ $6.762$ $6.734$ $6.221$ $6.115$ $5.987$ $5.985$ $6.247$ $6.246$ $6.997$ $6.995$ $7.498$ $7.497$ - -
Cinc 2019 ($10^{1}$) $7.456$ $6.597$ $7.347$ $7.311$ $7.212$ $7.246$ $5.889$ $6.072$ $7.608$ $7.541$ $10.293$ $9.7724$ $9.591$ $8.322$ - -
Heart Rate ($10^{-1}$) $1.195$ $1.001$ $1.123$ $0.950$ $0.986$ $0.845$ $0.774$ $0.633$ $0.627$ $0.497$ $0.945$ $0.683$ $0.671$ $0.587$ $1.753$ $0.832$
MIT-BIH $0.293$ $0.198$ $0.199$ $0.141$ $0.173$ $0.110$ $0.453$ $0.354$ $0.705$ $0.627$ $0.882$ $0.623$ $0.242$ $0.196$ $0.335$ $0.141$
CDC-IHA ($10^{1}$) $5.976$ $4.684$ $5.729$ $4.502$ $5.534$ $4.401$ $5.211$ $4.103$ $9.522$ $7.974$ $10.068$ $9.052$ $7.892$ $6.766$ $15.633$ $4.408$
@@ 4.2 Performance on Out-of-distribution Forecasting
Results. Detailed results of out-of-distribution performance are reported in Table 2. First, MIRA consistently achieves state-of-the-art performance, outperforming all general-domain foundation models and specialized time series baselines. Specifically, MIRAlarge achieves the best results on all datasets, improving RMSE by over 8% on average compared to the strongest baselines, confirming the advantage of scaling and medical pretraining. This advantage is particularly pronounced on clinical benchmarks such as MIT-BIH and CDC-IHA, where MIRAlarge achieves both the lowest RMSE and MAE. Second, domain-specific pretrnaining proves essential. All model variants—continue pretrained on medical corpora—consistently outperform models trained on general time series data. This demonstrates the benefit of leveraging medical-specific temporal structures and distributions during pretraining. Interestingly, even smaller MIRA models surpass larger general-domain models, suggesting that data relevance is more critical than model size alone. Third, MIRA achieves performance close to or even exceeding fine-tuned full-shot models in several cases. For instance, $MIRA_{large}$ outperform than full-shot baseline on MIT-BIH and slight lower in CDC-IHA.
@@ 4.3 Performance on In-distribution Forecasting
Ours
MIRAsmall $3.15$ $2.82$ $3.28$ $2.95$ $3.41$ $3.08$ $3.54$ $3.20$ $3.67$ $3.34$ $3.80$ $3.47$ $3.93$ $3.61$ $4.07$ $3.75$ $4.22$ $3.90$ 0 0
MIRAbase $2.98$ $2.68$ $3.11$ $2.81$ $3.25$ $2.94$ $3.38$ $3.07$ $3.51$ $3.20$ $3.64$ $3.34$ $3.78$ $3.48$ $3.93$ $3.63$ $4.10$ $3.79$ 0 11
MIRAlarge $2.85$ $2.56$ $2.95$ $2.65$ $3.08$ $2.78$ $3.21$ $2.90$ $3.35$ $3.04$ $3.49$ $3.18$ $3.63$ $3.31$ $3.99$ $3.86$ $3.97$ $3.64$ 88 88
General Setting
Time-MoEbase $3.05$ $2.78$ $3.18$ $2.92$ $3.32$ $3.05$ $3.45$ $3.18$ $3.58$ $3.31$ $3.71$ $3.43$ $3.85$ $3.57$ $3.99$ $3.71$ $4.15$ $3.86$ 0 0
Time-MoElarge $3.20$ $2.91$ $3.33$ $3.04$ $3.47$ $3.17$ $3.60$ $3.29$ $3.73$ $3.41$ $3.86$ $3.54$ $3.99$ $3.67$ $4.13$ $3.81$ $4.28$ $3.96$ 0 0
Moirailarge $4.80$ $4.18$ $5.39$ $4.64$ $5.68$ $4.55$ $6.02$ $4.83$ $6.09$ $5.06$ $6.18$ $5.02$ $6.31$ $5.09$ $6.51$ $5.24$ $6.71$ $5.36$ 0 0
Moiraibase $4.84$ $4.15$ $4.93$ $4.10$ $5.20$ $4.23$ $5.64$ $4.53$ $5.78$ $4.61$ $5.80$ $4.60$ $5.95$ $4.97$ $5.95$ $4.73$ $6.06$ $4.87$ 0 0
Moiraismall $4.91$ $4.22$ $5.04$ $4.20$ $6.12$ $4.93$ $6.23$ $4.98$ $6.44$ $5.49$ $6.47$ $5.18$ $6.59$ $5.60$ $6.66$ $5.20$ $6.86$ $5.71$ 0 0
Chronoslarge $4.71$ $4.10$ $4.85$ $4.22$ $5.02$ $4.31$ $5.31$ $4.52$ $5.58$ $4.70$ $5.82$ $4.91$ $5.97$ $5.10$ $6.15$ $5.32$ $6.43$ $5.49$ 0 0
Chronosbase $5.02$ $4.40$ $5.21$ $4.51$ $5.43$ $4.68$ $5.62$ $4.81$ $5.85$ $4.93$ $6.02$ $5.12$ $6.21$ $5.31$ $6.40$ $5.49$ $6.67$ $5.62$ 0 0
Chronossmall $5.32$ $4.71$ $5.49$ $4.82$ $5.68$ $4.93$ $5.81$ $5.02$ $6.01$ $5.20$ $6.21$ $5.32$ $6.40$ $5.49$ $6.58$ $5.61$ $6.83$ $5.72$ 0 0
Medical Setting
Time-MoElarge $2.95$ $2.70$ $3.08$ $2.83$ $3.22$ $2.96$ $3.35$ $3.09$ $3.48$ $3.22$ $3.61$ $3.35$ $3.75$ $3.49$ $3.89$ $3.63$ $4.04$ $3.78$ 11 0
Time-MoEbase $3.05$ $2.78$ $3.18$ $2.91$ $3.32$ $3.04$ $3.45$ $3.17$ $3.58$ $3.30$ $3.71$ $3.43$ $3.85$ $3.57$ $3.99$ $3.71$ $4.15$ $3.86$ 0 0
Moirailarge $4.41$ $3.76$ $4.92$ $4.20$ $5.35$ $4.41$ $5.69$ $4.76$ $5.93$ $4.80$ $6.09$ $4.92$ $6.23$ $4.98$ $6.48$ $5.29$ $6.62$ $5.30$ 0 0
Moiraibase $4.55$ $3.92$ $4.78$ $4.07$ $5.18$ $4.26$ $5.49$ $4.46$ $5.78$ $4.64$ $5.70$ $4.51$ $5.87$ $4.97$ $5.92$ $4.75$ $5.96$ $4.71$ 0 0
Moiraismall $5.02$ $4.21$ $5.36$ $4.40$ $5.55$ $4.98$ $5.97$ $4.92$ $6.25$ $5.34$ $6.43$ $5.11$ $6.44$ $5.35$ $6.57$ $5.41$ $6.58$ $5.30$ 0 0
Chronoslarge $4.30$ $3.89$ $4.58$ $4.02$ $4.83$ $4.20$ $5.02$ $4.31$ $5.32$ $4.51$ $5.61$ $4.73$ $5.83$ $4.91$ $6.02$ $5.10$ $6.31$ $5.32$ 0 0
Chronosbase $4.61$ $4.10$ $4.83$ $4.20$ $5.02$ $4.31$ $5.32$ $4.51$ $5.58$ $4.70$ $5.82$ $4.91$ $5.97$ $5.10$ $6.15$ $5.32$ $6.43$ $5.49$ 0 0
Chronossmall $4.83$ $4.31$ $5.02$ $4.43$ $5.32$ $4.51$ $5.58$ $4.70$ $5.82$ $4.91$ $5.97$ $5.10$ $6.15$ $5.32$ $6.43$ $5.49$ $6.71$ $5.61$ 0 0